Okay it says to solve by completeing the square.
4x^2+2x-3=0
So far I have the following:
4x^2+2x=3
x^2+2/4x=3/4
Then I get totally lost and don't know what else to do!! HELP! PLEASE!
Thanks in advance!!!!!!
CH

Found 2 solutions by venugopalramana, jai_kos:
Answer by venugopalramana(3286)

Okay it says to solve by completeing the square.
4x^2+2x-3=0
So far I have the following:
4x^2+2x=3
x^2+2/4x=3/4
X^2+X/2-3/4=0..........MAKE PERFECT SQUARE USING X^2 AND X TERMS WITH THE FORMULA (X+A)^2 = X^2+2XA+A^2...THAT IS 2XA=X/2.....2A=1/2...A=1/4....
NOW ADD AND SUBTRACT A^2...THAT IS (1/4)^2....
=[(X)^2 + 2(X)(1/4)+(1/4)^2]-(1/4)^2-3/4=0
= [X + (1/4)]^2= 1/16 +3/4 = 13/16
X+1/4 = + OR - SQRT(13)/4
X = -1/4 + OR - SQRT(13)/4 = [-1 + OR - SQRT(13)]/4

Then I get totally lost and don't know what else to do!! HELP! PLEASE!
Thanks in advance!!!!!!

Answer by jai_kos(139)

Given 4x^2+ 2x -3 = 0
Get all the “x” terms to one side and terms on the other side.
We get
4x^2 + 2x = 3
Let us make the above equation even.
Multiply the above equation by 4.
16x^2 +8x = 12
Add one on both sides.
16x^2 + 8x + 1 = 12 +1
16x^2 + 4x + 4x + 1 = 13
4x( 4x +1) +1(4x +1) = 13
(4x +1)^2 = 13
Taking square root on both sides, we get
4x +1 = + Sqrt(13)
Subtract 1on both sides, we get
4x = + Sqrt(13) -1
Rearrange the above, we get
4x = -1 + Sqrt(13)
Divide by 4 in the above equation, we get
x = [-1 + Sqrt(13)] / 4